Jack is now seven times as old as James. In 5 years, Jack will be five times as old as James. How old is James now?
70
55
10

Answer by VFBundy(438)

James's age = x
Jack's age = y

y = 7x
y + 5 = 5(x + 5)

y = 7x
y + 5 = 5x + 25

7x - y = 0
5x - y = -20

7x - y = 0
-5x + y = 20

2x = 20

x = 10

James's age = x = 10